I am horrible at math. Can you please help solve.
Suppose that N= cuberoot x plus 2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?

Answer:

Show your work here:
b) In approximately how many years will there be 7 million cases?
Answer:

Show your work here:
: I am horrible at math. Can you please help solve.
Suppose that N= cuberoot x plus 2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?

Answer:

Show your work here:
b) In approximately how many years will there be 7 million cases?
Answer:

Show your work here:

Answer by ankor@dixie-net.com(4484)

Suppose that N= cuberoot x plus 2 models the number of cases of an infection, in millions, of a disease x years from now.
:
WE can write the equation like this:
N = x^(1/3)+ 2
a) Approximately how many cases of the infection will there be 16 years from now?
N = 16^(1/3)+ 2
N = 2.5198 + 2; found the cube root of 16 on a calc
N = 4.52 million in 16 yrs
:
:
b) In approximately how many years will there be 7 million cases?
x^(1/3)+ 2 = 7
Subtract 2 from both sides:
x^(1/3) = 7 - 2
x^(1/3) = 5
Cube both sides and you have;
x = 5^3
x = 125 yrs
:
:
It occurred to me that you may have meant the formula to be:
N = (x+2)^(1/3); you can let me know if that is the case and I will edit the procedure.